OpenStudy (anonymous):

passes through (-2,0) and is perpendicular to the line whose equation is y=-3x+7

hero (hero):

Okay, going to go watch Smallville now...

OpenStudy (anonymous):

smallvile is over

OpenStudy (anonymous):

ok?

hero (hero):

Not for me it isn't

OpenStudy (anonymous):

hahahaha

hero (hero):

I'm only on Season 4

OpenStudy (anonymous):

you wathching it online

hero (hero):

Yes, through Amazon

OpenStudy (anonymous):

hi guys... um i kinda need help.

hero (hero):

Ooops

OpenStudy (anonymous):

cool, it was a good show

hero (hero):

Heard it got kind of lame toward the end

OpenStudy (anonymous):

its just that i would have liked clark to become like the full blown superman

OpenStudy (ybarrap):

y=-3x+7 yperp = (1/3)x + b yperp(-2) = 0 = (1/3)(-2) + b b = 2/3 Ans. yperp = (1/3)x + 2/3

OpenStudy (anonymous):

the show just is kind of about his life, personal life

OpenStudy (anonymous):

it should be -1/3x -2/3 i think

hero (hero):

They'll come out with a Superman movie of course

OpenStudy (anonymous):

what does yperp mean?

OpenStudy (anonymous):

y perpendicular

OpenStudy (anonymous):

oh right.... its too late for this stuff. my brain cant process!

OpenStudy (anonymous):

the equation is y= -1/3X - 1/3 actually haha, those guys were random

OpenStudy (anonymous):

yeah ikr... like um hi! nice to meet you!

OpenStudy (anonymous):

haha, oh, i didn't see that it was y=-3x+7, mahhh bad, he was right

OpenStudy (anonymous):

yeah thats what it was (: i have no clue what im doing. i got assigned a summer packet and i never did it, so now im stuck doing it. ive never even learned this stuff before. im like on my own!

OpenStudy (anonymous):

but i just have one more question if you have time.....

OpenStudy (anonymous):

sorry, i had to get something to eat, you still need help?